OpenStudy (anonymous):

Find the values of a so that the equation (a^2-1)x^2+2(a-1)x+1>0

OpenStudy (turingtest):

not sure what to do here frankly

OpenStudy (anonymous):

Should be finding domain or something.

OpenStudy (zarkon):

you could complete the square

OpenStudy (turingtest):

agreene and I did that on the other post, but then what do we do with the result?

OpenStudy (anonymous):

y=(a^2-1)x^2+2(a-1)x+1 is the original task. Then it says to find the values of a so that the equation would be positive. Answer should be a=(1;infinity)

OpenStudy (agreene):

\[a\ge 1\] there are a few other--rather odd things you can do the get the rest of the answers.

OpenStudy (anonymous):

And how did you get it?

OpenStudy (turingtest):

yeah I see now agreen showed that this the expression for x in terms of a is not defined for x<1 look at the other post

OpenStudy (agreene):

the rest of the solutions... are just moving around the answer from the x=(junk) http://www.wolframalpha.com/input/?i=solve+%28a%5E2-1%29x%5E2%2B2%28a-1%29x%2B1%3E0

OpenStudy (anonymous):

One possible answer \( -2<x\leq 0 \)

OpenStudy (turingtest):

\[x=\frac{\pm i\sqrt{2}-\sqrt{a-1}}{\sqrt{a-1}(a+1)}\]so this is imagniary when a<1 does that pose a problem though? we already have an imaginary number up top....

OpenStudy (anonymous):

I am really not allowed to use imaginary numbers

OpenStudy (turingtest):

I didn't check agreen's work, so if that i popped up spuriously you have to ask him about it lol

OpenStudy (zarkon):

\[(a^2-1)x^2+2(a-1)x+1=(a^2-1)\left(x+\frac{1}{a+1}\right)^2+\frac{2}{a+1}\]

OpenStudy (agreene):

it causes issues for when a<1... in the sense that you end up with answers that are not legitimate--which is why the other answers included x is an element of the reals so, we know to ignore trivial answers. but a >= 1 is a pretty solid answer, the others are very messy and probably extraneous to what ErkoT is doing.

OpenStudy (anonymous):

And I get the answer by doing?

OpenStudy (turingtest):

I'm still trying to figure out where Zarkon is going with his factoring out of a^2-1

OpenStudy (zarkon):

\[(a^2-1)\left(x+\frac{1}{a+1}\right)^2+\frac{2}{a+1}\] if \(a\ge1\) then the number is clearly positive if \(-1<a<1\) then \(a^2-1\) is negative so choose x large enough so that \((a^2-1)\left(x+\frac{1}{a+1}\right)^2\) dominate \(\frac{2}{a+1}\) if \(a\le 1\) choose \(x=1/(a+1)\) then the function is negative. thus only\[[1,\infty)\] works

OpenStudy (zarkon):

typo if a<-11 choose x=-1/(a+1) then the function overall is negative.

OpenStudy (zarkon):

geesh...a<-1

OpenStudy (zarkon):

not sure why your original answer does not include 1

OpenStudy (anonymous):

I have absolutely no idea.

OpenStudy (turingtest):

I'm gonna process that answer while I go enjoy my Sunday playing Skyrim :P

OpenStudy (zarkon):

\[(1^2-1)x^2+2(1-1)x+1=1>0\]

OpenStudy (agreene):

I'm gonna play some Starwars the Old Republic to clear my mind i think >.>

OpenStudy (zarkon):

any MW3 players :)

OpenStudy (agreene):

I used to play MW3... havent recently though

OpenStudy (anonymous):

I just played it an hour ago.

OpenStudy (turingtest):

not yet still got MW2 and BO everything comes late to Mexico but man, I wanna play you guys!

OpenStudy (zarkon):

you guys play on steam?

OpenStudy (turingtest):

steam?

OpenStudy (zarkon):

I guess not ;)

OpenStudy (zarkon):

http://store.steampowered.com/

OpenStudy (zarkon):

I play all of my games through steam

OpenStudy (turingtest):

Thanks, I'll check it out! PS3 games are so expensive in Mexico, everyone buys pirate Xbox games instead

OpenStudy (zarkon):

ah...this is a PC/Mac thing

OpenStudy (zarkon):

I don't play console games...only PC

OpenStudy (turingtest):

I'll switch to PC games for the next gen probably

OpenStudy (agreene):

Mexico... I've been there. Me encanta Mérida, Yucatán.

OpenStudy (anonymous):

Cool, you guys are from another continent.

OpenStudy (agreene):

Intercontinental Math Help!

OpenStudy (turingtest):

nuca he ido alla (no tengo acentos) vivo en Guadalajara visite a Mazamitla y Guanjuato but no, I'm actually american plus, mexico is the smae conctinent ;)

OpenStudy (turingtest):

nunca

OpenStudy (zarkon):

ErkoT: where are you

OpenStudy (agreene):

never went to Guadalajara, meant to... but just didnt.

OpenStudy (anonymous):

Europe.

OpenStudy (zarkon):

I've been to Germany...but that is it

OpenStudy (turingtest):

I'd like to know the demographics on this site lots of Middle-east as well

OpenStudy (agreene):

Yeah, lots of Asians--Indian it seems

OpenStudy (agreene):

I havent made it to Europe yet... probably will after I finish my Masters Degree.

OpenStudy (turingtest):

me too

OpenStudy (zarkon):

I was there in 1993 and 1996

OpenStudy (turingtest):

learn any German? enough to read all those cool German math papers?

OpenStudy (zarkon):

no

OpenStudy (turingtest):

lol

OpenStudy (agreene):

I was 6 and 9 years old, respectively during that time. Ich liebe Deutsch

OpenStudy (anonymous):

In 1993 I wasn't even born yet.

OpenStudy (agreene):

sshh you

OpenStudy (turingtest):

right!

OpenStudy (zarkon):

I was in the military...it was for training purposes

OpenStudy (turingtest):

whoa, didn't expect that o-0

OpenStudy (anonymous):

Same.

OpenStudy (zarkon):

I don't have your typical mathematician background :)

OpenStudy (anonymous):

I don't have it either.

OpenStudy (turingtest):

Nor I You are definitely a mysterious figure Zarkon, and now even more so

OpenStudy (zarkon):

good...I like it that way ;)

OpenStudy (turingtest):

figured as much ;)

OpenStudy (anonymous):

Alright guys, I'm off. Thank you for your help!